OpenStudy (anonymous):

HELP I WILL FAN AND MEDAL @kelliegirl33

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

there is 5 questions

OpenStudy (anonymous):

I'm here...whats up lala

OpenStudy (anonymous):

What is the length of the line segment that is graphed from (+2, +5) to (+8, +5)? A. 3 units B. 4 units C. 6 units D. 10 units

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

distance formula = sq rt (x2 - x1)^2 + (y2 - y1)^2 (2,5) x1 = 2 and y1 = 5 (8,5) x2 = 8 and y2 = 5 now we sub d = sq rt (8 - 2)^2 + (5 - 5)^2 d = sq rt (6^2) + (0^2) d = sq rt 36 d = 6 distance is 6 units

OpenStudy (anonymous):

an easier way, I just noticed, is ....well you notice that your y's have the same number. That means you line is horizontal. You can just take the distance of your x points....2 to 8 = 6 units

OpenStudy (anonymous):

that was so rude what lisa said to you...she had no right

OpenStudy (anonymous):

A line segment is drawn from (+1, +3) to (+8, +3) on a coordinate grid. Which answer explains one way that the length of this line segment can be determined? A. Subtract 8 – 1. B. Subtract 8 – 3. C. Add 1 + 3. D. Add 1 + 8.

OpenStudy (anonymous):

subtract 8 - 1

OpenStudy (anonymous):

What is the length of the line segment shown on this coordinate grid? A. 2 units B. 3 units C. 4 units D. 6 units

OpenStudy (anonymous):

need the line segmants

OpenStudy (anonymous):

you have a vertical line...count how many spaces are between the points and what do you get ?

OpenStudy (anonymous):

one point is on 6 and 1 point is on 2....whats 6 - 2 ?

OpenStudy (anonymous):

or how many spaces are between 6 and 2

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

when you have a vertical line (both x's are the same) or a horizontal line (both y's are the same), then you can take the distance of the other variables by subtracting one from the other. The answer to this question is 4 units...6 - 2 = 4 understand ?

OpenStudy (anonymous):

its just a matter of counting when it comes to vertical or horizontal lines. However, if they are not vertical or horizontal, you would use the formula I used in the first problem

OpenStudy (anonymous):

A line segment is drawn from (+4, +2) to (+4, +5) on a coordinate grid. Which answer explains one way that the length of this line segment can be determined? A. Subtract 4 – 2 B. Subtract 5 – 2 C. Add 4 + 2 D. Add 2 + 5

OpenStudy (anonymous):

same x points....so what is the difference in the y points...the difference is: subtract 5 - 2

OpenStudy (anonymous):

Which answer explains one way that the length of this line segment can be determined? A. Add 4 + 7. B. Add 1 + 4. C. Subtract 7 – 4. D. Subtract 4 – 1.

OpenStudy (anonymous):

subtract 7 - 4

OpenStudy (anonymous):

do you have more lala ?